OpenStudy (faman39):

The label on a 2-liter bottle of soda indicates that one serving is 8 fl oz and that there are 8 servings in the bottle. Are these exact measurements? Explain, using mathematics to justify your answer.

OpenStudy (kymber):

First you need to find out how many ounces 2 liters is

OpenStudy (kymber):

HELLO? YOU KNOW WHAT IT IS?

OpenStudy (anonymous):

The Formulas for conversion from Fahrenheit to Celsius and vice versa are as follows: C = 5/9 * (F - 32) F = 9/5C + 32 Water boils at 212°F which is 100°C, so plug those numbers into the formula to check it: 100 = 5/9 * (212-32). 5/9 of 180 = 100 212 = 9/5(100) + 32. 9/5 of 100 + 32 = 212. As for the 2-liter bottle problem, the 8 servings/bottle 8 fl oz cannot be exact measurements because 1 Liter is equivalent to 33.814 ounces. Therefore, a 2-liter bottle has 67.628 ounces, and if we divide this by 8 servings we get: 8.45 oz - which is a bit more than what the label claims.